Find the greatest of three consecutive odd integers if the product of the second and third is 8 more than 13 times the first.
Answer by stanbon(75887)

Find the greatest of three consecutive odd integers if the product of the second and third is 8 more than 13 times the first.
---
1st: 2x-3
2nd: 2x-1
3rd: 2x+1
----------------------
Equation:
(2x-1)(2x+1) = 13(2x-3)+8
---
4x^2-1 = 26x-31
---
4x^2-26x+30 = 0
2x^2-13x+15 = 0
---
2x^2-10x-3x+15 = 0
2x(x-5)-3(x-5) = 0
(x-5)(2x-3) = 0
x = 5 or x = 3/2
====================
If x = 5:
2x-3 = 7
2x-1 = 9
2x+1 = 11
---
If x = 3/2
2x-3 = 0
2x-1 = 2
2x+1 = 4
These are not odd, so x = 3/2 is not a solution.
============================
Final answer: 2x+1 =11
=============================
Cheers,
Stan H.
Cheers,
Stan H.
====================